OpenStudy (anonymous):

What is the derivative of (250sqrt(x) + ((x^2)/15625))/x ?

OpenStudy (anonymous):

i would like to help , just gemme time k

OpenStudy (anonymous):

k thank you

OpenStudy (anonymous):

how is ((x^2)/15625))/x ?

OpenStudy (anonymous):

well i don't know about this , sorry

OpenStudy (kropot72):

Is this the expression to be differentiated? \[\frac{250\sqrt{x}+\frac{x ^{2}}{15625}}{x}\]

OpenStudy (anonymous):

I find it funny that I actually have more trouble doing differentiations than integrations. You can give me a trigonometric fraction and I could (probably) integrate it just fine. But give me a problem like this and my eyes will dilate and I will cower in fear. But on a more serious note, I believe you should start this problem out by splitting the fraction into two fractions. Since you've got two terms on top divided by one term on the bottom, you can easily split the 250*sqrtx as one fraction (divided by x) and then the x^2/(15625) as the second fraction (divided by x again) and then adding the derivatives of the two fractions to get the final answer. BUT, before you take my word for it, PLEASE verify that statement with someone else; I might be making up rules because I KNOW that'd work for integrating, but I could be wrong by trying to apply it to differentiating.